# Scattering theory of higher order topological phases

### v1

@@ Abstract
The surface states of intrinsic higher order topological phases are protected by the spatial symmetries of a finite sample. This property makes the existing scattering theory of topological invariants inapplicable because the scattering geometry is either incompatible with the symmetry or does not probe the bulk topology. We resolve this obstacle by using a symmetric scattering geometry that probes transport from the inside to the outside of the sample. We demonstrate that the intrinsic higher order topology is captured by the flux dependence of the reflection matrix. Our finding follows from identifying the spectral flow of a flux line as a signature of higher order topology. We show how this scattering approach applies to several examples of higher order topological insulators and superconductors. Our theory provides an alternative approach for proving bulk–edge correspondence in intrinsic higher order topological phases, especially in presence of disorder.
@@ 1 Introduction
Despite the different nature of the surface states, important properties of topological insulators also apply to higher order phases. Firstly, there are higher order phases that are stable in the presence of disorder, similar to the strong topological phases [6, 7, 8, 9]. This protection follows from the observation that removing a protected surface state requires coupling it with its symmetry partner, which, however, is located at the opposite edge of the sample. Therefore, because disorder acts locally, it may not remove the protected surface states. Secondly, several works used flux response and dislocation defects as a signature of higher order topology [10, 11, 12, 13, 14, 15, 16, 17], generalizing the Laughlin argument. This is reminiscent of the spectral flow in strong topological phases. We thus identify two open questions:
@@ 2 Why the extrinsic scattering invariant is not enough
The most direct approach to applying a scattering invariant to a HOTI is to use the appearance of the protected zero modes on the edge. This requires probing a reflection matrix of a finite symmetric sample with leads attached to its edges in a symmetric way. Symmetry class–dependent functions of the reflection matrix at each lead count the number of protected zero modes at the edge, and their parity was used as a scattering probe for HOTIs [3]. This extrinsic scattering invariant, and similar real space HOTI invariants [25, 26], rely on knowing the position of protected edge states, and therefore its usefulness in characterising an unknown system is limited.
@@ 3 Intrinsic scattering invariant construction
where $H$ is the Hamiltonian of the scattering region and the leads, and $\Psi^{\textrm{in}}$, $\Psi^{\textrm{out}}$, and $\Psi^{\textrm{localized}}$ are the wavefunctions of the incoming, outgoing, and localized states, respectively. These wavefunctions are matrices with columns corresponding to different modes, see App. A for a detailed description. To treat topology due to antisymmetries that map $E$ to $-E$ on the same footing, we set the energy $E$ to zero. The incoming and outgoing wavefunctions occupy the same lead sites and orbitals, while the localized wavefunctions are only defined in the scattering region. The total number of modes is the number of columns of the wavefunction matrices $N_{\textrm{modes}}=N_{\textrm{sites}}\times N_{\textrm{orbitals}}$, where $N_{\textrm{sites}}$ is the number of sites to which the leads are attached, and $N_{\textrm{orbitals}}$ is the number of orbitals per site in the lead. The scattering equations define a valid solution for any combination of amplitudes $\alpha^{\textrm{in}}$ of the incoming modes. Because the scattering matrix in Eq. (1) specifies a relation between the amplitudes of the incoming wavefunctions to those of the outgoing ones, the scattering matrix is a linear map rather than an operator. Separating the leads into two groups imposes a block structure on the scattering matrix:
@@
Figure 3: Scattering geometries that probe the existence of zero-energy surface states using the bulk symmetries. The reflection matrix $r$ encodes the amplitudes of the wavefunctions that are reflected back into the lead (purple) from the scattering region (turquoise). (a) Scattering setup used to probe topological phases protected by local symmetries. A two-dimensional bulk Hamiltonian defines a one-dimensional reflection matrix by attaching leads along one direction and applying twisted boundary conditions in the other. The reflection matrices $r$ and $r^{\prime}$ capture the appearance of zero-energy bound states (b). (c) Scattering setup used to probe topological phases protected by spatial symmetries. A two-dimensional six-fold symmetric bulk Hamiltonian defines a one-dimensional reflection matrix by attaching symmetric leads and introducing a flux through the center of the scattering region. The blue area shows the matrix elements of the symmetry operator that gain a phase factor due to the flux, defining $C_{6}(\phi)$. The scattering setup probes the reflection from the leads attached to the outer ring, all of which are considered as one lead.
The scattering theory of topological invariants [19] relates the topology of a $d$-dimensional bulk Hamiltonian $H_{\textrm{bulk}}(\bm{k}_{d})$, with $\bm{k}_{d}$ the wave vector, to the topology of a $d-1$-dimensional effective Hamiltonian $H_{\textrm{eff}}(\bm{k}_{d-1})$. The dimensional reduction procedure of obtaining this effective Hamiltonian is as follows:
@@
    $H_{\textrm{eff}}=\begin{cases}r&\text{if }H_{\textrm{bulk}}\text{ has chiral symmetry,}\\ \begin{pmatrix}0&r\\ r^{\dagger}&0\end{pmatrix}&\text{otherwise}.\end{cases}$ (3)
This effective Hamiltonian has zero eigenvalues simultaneously with $r(\bm{k}_{d-1})$ and a symmetry class different from the symmetry of $H_{\textrm{bulk}}$. The zero eigenvalues of $r(\bm{k}_{d-1})$ correspond to quantized transmission eigenvalues from one lead to the other, which may only happen if the gap of $H_{\textrm{bulk}}$ closes. Because of this relation, the topological invariants of $H_{\textrm{eff}}$ and $H_{\textrm{bulk}}$ are equal. An alternative interpretation of the scattering invariant is to consider a boundary between the topological bulk and a trivial region, so that the boundary Green’s function defines the topology of the bulk [30]. In the scattering description this boundary corresponds to interrupting the lead with a reflection matrix $r^{\prime}$ of the trivial region, as shown in Fig. 3(b). The zero energy modes at an interface of an infinite system appear whenever $rr^{\prime}$ has an eigenvalue equal to $1$, and therefore the protected zero energy solutions at the interface are encoded in $r$.
While the procedure above may be applied to intrinsic HOTIs, the resulting scattering geometry either has leads swapped by the symmetry of the HOTI, or is incompatible with the global spatial symmetry protecting the HOTI phase. This incompatibility makes the topology of $H_{\textrm{eff}}$ different from the higher order topology of $H_{\textrm{bulk}}$. Instead, in order to apply the scattering approach to an intrinsic HOTI, the scattering geometry must be compatible with the global spatial symmetry protecting the HOTI phase, and this symmetry must map the leads onto themselves. At the same time, the two leads must be separated by the bulk rather than belonging to the same surface. This is required in order for zeros of the reflection matrix to correspond to the bulk gap closing and therefore to probe only bulk topological transitions. Otherwise, zeros of the reflection matrix could appear due to the closing of the surface gap and therefore not probe the bulk topology, as illustrated in Sec. 2. To satisfy the requirements of the scattering theory we propose to use a finite geometry with a hole: an annulus in two dimensions or a cylinder in three dimensions, and attach one lead to the inside and one lead to the outside of the sample, as shown in Fig. 3(c).
@@
Similar to the dimensional reduction procedure, we apply twisted boundary conditions in the direction perpendicular to the cylinder axis, which introduces a wave vector $k_{d-2}$. Without introducing additional parameters, this would result in a $d-2$-dimensional reflection matrix, which loses information about the angular direction along the cylinder, and therefore is unlikely sufficient to determine the HOTI topology. To preserve the information about the angular direction, we therefore introduce a flux $\phi$ through the center of the cylinder, which acts as an additional momentum parameter. The resulting procedure maps a $d$-dimensional bulk onto a $d-1$-dimensional $H_{\textrm{eff}}(\phi,\bm{k}_{d-2})$, similar to the dimensional reduction of the strong topological insulators. Contracting the inner radius of the cylinder to a point generally gaps out the surface modes due to the finite size effects, and therefore the flux is necessary to cancel this splitting, as shown in Fig. 4. This cancellation of the finite size splitting was reported in Ref. [31], in a specific symmetry class. By confirming that the scattering invariants work in the annulus geometry, we show that the flux response is a general feature of the higher order topology.
@@
where $\tilde{H}(R_{\mathcal{O}}(\phi,\bm{k}_{d-2}))=\mathcal{O}H(\phi,\bm{k}_{d-2})\mathcal{O}^{-1}$ is the transformed Hamiltonian, and $R_{\mathcal{O}}$ is the action of the operator in parameter space. To obtain the transformed scattering matrix $\tilde{S}(R_{\mathcal{O}}(\phi,\bm{k}_{d-2}))$, we act with the operator on $\mathcal{O}$ on the wavefunctions and coefficients of the wavefunctions, and equate the coefficients of the wavefunctions in Eq. (1) and Eq. (4) (see App. B for a detailed derivation). The action of an operator on the lead wavefunctions depends on whether this operator is unitary or antiunitary, and whether it is a symmetry or an antisymmetry of the Hamiltonian. We denote the unitary symmetries $\mathcal{U}$, unitary antisymmetries $\mathcal{C}$ (chiral symmetry), antiunitary symmetries $\mathcal{T}$ (time-reversal symmetry), and antiunitary antisymmetries $\mathcal{P}$ (particle-hole symmetry). The action of these 4 types of operators on the lead modes is [19] [^2] Note that we choose a different convention for $V$ and $Q$ than in Ref. [19]. Swapping $V\leftrightarrow V^{T}$ and $Q\leftrightarrow Q^{T}$ would make our convention equivalent to Ref. [19].:
@@
where $V_{\mathcal{O}}(\phi,\bm{k}_{d-2})$ and $Q_{\mathcal{O}}(\phi,\bm{k}_{d-2})$ are unitary matrices, and which we compute by projecting the operator $\mathcal{O}$ onto the lead wavefunctions. Therefore, combining the transformed scattering equation Eq. (4) with the transformed wavefunctions in Eqs. (5), we obtain the transformed reflection matrix $\tilde{r}$:
@@
Due to $r(\phi,\bm{k}_{d-2})$ being a linear map (as a submatrix of $S$), the symmetry operations (6) apply to $r(\phi,\bm{k}_{d-2})$ differently than to $H_{\textrm{bulk}}$. Additionally, due to Eq. (3), whenever $H_{\textrm{bulk}}$ has chiral symmetry, $H_{\textrm{eff}}$ does not, and vice versa. This results in Altland–Zirnbauer symmetry class of $H_{\textrm{bulk}}$ being related to that of $H_{\textrm{eff}}$ according to the Bott periodicity [19]. For example, in the symmetry class D ($\mathcal{P}^{2}=1$), particle-hole symmetry constraint is equivalent to $H_{\textrm{bulk}}(\mathbf{k})=-H_{\textrm{bulk}}^{*}(-\mathbf{k})$ up to a basis choice. The reflection matrix symmetry constraint (6b) is then equivalent to $r(-\phi,\bm{k}_{d-2})=r^{*}(-\phi,-\mathbf{k}_{d-2})$. Then, after applying (3), the effective Hamiltonian then belongs to the symmetry class BDI ($\mathcal{T}^{2}=\mathcal{P}^{2}=1$).
@@
In practice, the first step of determining the topological invariant requires finding a basis for the incoming and outgoing wavefunctions where the symmetry constraints on the reflection matrix are minimal. Choosing an appropriate basis simplifies the time-reversal symmetry constraint to either symmetry or antisymmetry of $r(\phi,\bm{k}_{d-2})$, particle-hole symmetry becomes a reality constraint, and chiral symmetry becomes a Hermiticity constraint, as shown in App. B. In the presence of multiple symmetries, we choose the basis where the chiral symmetry gives a Hermiticity constraint if it is present, according to Eq. (3), and the other symmetries have a minimal compatible form, following Ref. [19]. After performing this transformation, we rely on the established theory of Hamiltonian topological invariant to determine the topological invariant of $H_{\textrm{eff}}$.
@@ 4.1 Intrinsic scattering invariant of BBH model
To construct the intrinsic scattering invariant of the BBH model, we first consider the reflection matrix of an annulus-shaped lattice with a flux $\phi$, as shown in Fig. 5(a). Four-fold rotation and sublattice symmetries given by Eqs. (6a) and (6d) constrain the reflection matrix $r(\phi)$ as
$r(\phi)=Q_{C_{4}}(\phi)\>r(\phi)V_{C_{4}}^{\dagger}(\phi)=V_{\mathcal{C}}\>r(\phi)^{\dagger}Q_{\mathcal{C}}^{\dagger},$ (10)
@@
As the first step, we simplify the problem by observing that $C_{2}\equiv C_{4}^{2}$ commutes both with $\mathcal{C}$ and $\mathcal{C}_{4}$. Because $C_{2}^{2}(\phi)=-\exp(i\phi)$, interpreting $\phi$ as a momentum parameter of the effective Hamiltonian means that the problem is invariant under translation by half a unit cell in the corresponding direction. Therefore, we unfold the Brillouin zone and reduce the unit cell to a primitive one to get rid of redundant degrees of freedom. Because the eigenvalues of $C_{2}$ are ±$\pm i\exp(i\phi/2)$, they are periodic over $4\pi$ and the two subspaces of $C_{2}$—and of any operator that commutes with it—swap over a $2\pi$ interval. Therefore, we project $r$ and the remaining symmetries onto the subspace of $C_{2}$ with the eigenvalue $i\exp i\phi/2$, and consider their dependence on $\phi\in[0,4\pi)$. The range of $\phi\in[2\pi,4\pi)$ of the $i\exp i\phi/2$ eigenspace is equivalent to the range $\phi\in[0,2\pi)$ of the $-i\exp i\phi/2$ eigenspace, so choosing one subspace over twice the interval is an equivalent representation. To avoid dealing with $4\pi$ periodicity, we redefine $\phi:=\phi/2$, which yields $r(\phi)$ that satisfies symmetry constraints equivalent to those in Eq. (10), but with $V_{C_{4}}^{2}=Q_{C_{4}}^{2}=i\exp i\phi$ instead. We call this procedure factoring out a symmetry, because $r(\phi)$ is no longer constrained by $C_{2}$. We describe how to factor out a symmetry and transform the remaining symmetries to the new basis in App. C.
@@
where the two blocks of $H_{\textrm{eff}}(\phi)$ correspond to different eigensubspaces of $V_{C_{4}}$. Here we use that the blocks of $V_{C_{4}}$ swap under $\phi\to\phi+2\pi$, and therefore the effective Hamiltonian in Equation (13) is $4\pi$-periodic in $\phi$. Contrary to the dimensional reduction procedure used for strong topological insulators, this Hamiltonian has a sublattice-like symmetry—more specifically, a glide sublattice symmetry—just like its parent BBH model. The final step is to apply the topological invariant of $H_{\textrm{eff}}(\phi)$, however, to the best of our knowledge, the topology of one-dimensional systems combining fractional translations and sublattice symmetries has not been studied before. Therefore, we proceed to construct the invariant, although in most cases one may apply the already established invariants to the effective Hamiltonian.
Because $H_{\textrm{eff}}(\phi)$ is gapped as long as $h(\phi)$ has no zero eigenvalues, its topological invariant must be a function of $\det h(\phi)$. Because $\det h(\phi+2\pi)=\det h^{\ast}(\phi)$, $\det h(\phi)$ crosses the real axis an odd number of times in the interval $\phi\in[0,2\pi)$. The $\mathbb{Z}_{2}$ invariant of this system is the parity of the number of the crossings by $\det h(\phi)$ of the positive real axis: because the crossings of the real axis may only appear and disappear in pairs, the parity of the number of crossings may only change if $\det h(\phi)=0$ for some $\phi$. Therefore we identify the scattering invariant of the BBH model to be:
@@
We construct the model and compute the scattering matrix using Kwant [32]. To demonstrate that the scattering invariant is an effective tool to study disordered systems, we modify the BBH model by adding normally-distributed disorder in the intra-cell hopping $\lambda$ that is symmetric under the global $C_{4}$ symmetry. Figure 5 shows that the invariant correctly changes at the phase transition of the BBH model, even when the leads are not in close proximity to the corner charges. Because scattering invariants rely on the unitarity of the reflection matrix, the phase transition point is sensitive to finite-size effects that may arise from the overlap of the wavefunctions between the inner and outer radius of the annulus. Figures 5(b-f) show the gap of the reflection matrix $\Delta=\min|\det r(\phi)|$ and the invariant $\mathcal{Q}$ as a function of $\lambda/\gamma$ for different outer radii $r$ of the annulus. In the thermodynamic limit of $r\to\infty$ and complete absence of disorder, the invariant changes sign at $\lambda=\gamma$, as expected from the bulk topological invariant of the BBH model. Regardless of the presence of disorder or the size of the annulus, the invariant is always quantized and changes sign only if the reflection matrix has a zero eigenvalue.
Figure 5: Two scattering geometries of the BBH model and their intrinsic invariant in the presence of disorder. (a/d) Annulus geometry with a flux $\phi$ threaded through the center and leads (purple) attached to the inner and outer rings. In (a) the leads are attached where the zero modes are located and in (d) the leads are in between the zero modes. The gap of the reflection matrix $\Delta=\min|\det r(\phi)|$ in (b/e) closes simultaneously with the sign change of the invariant $\mathcal{Q}$ in (c/f). The results are shown for different outer ring sizes $r=10,20,40$, and in (c/f) the values of $\mathcal{Q}$ are minimally shifted for clarity.

### v2

@@ Abstract
The surface states of intrinsic higher order topological phases are protected by the spatial symmetries of a finite sample. This property makes the existing scattering theory of topological invariants inapplicable: the scattering geometry is either incompatible with the symmetry or does not probe the bulk topology. We resolve this obstacle by using a symmetric scattering geometry that probes transport from the inside to the outside of the sample. We demonstrate that the intrinsic higher order topology is captured by the flux dependence of the reflection matrix. Our finding follows from identifying the spectral flow of a flux line as a signature of higher order topology. We show how this scattering approach applies to several examples of higher order topological insulators and superconductors. Our theory provides an alternative approach for proving bulk–edge correspondence in intrinsic higher order topological phases, especially in the presence of disorder.
@@ 1 Introduction
Despite the different nature of the surface states, important properties of topological insulators also apply to higher order phases. First, some higher order phases are stable in the presence of disorder, similar to strong topological phases [6, 7, 8, 9, 10]. This protection follows from the observation that removing a protected surface state requires coupling it with its symmetry partner, which, however, is located at the opposite edge of the sample. Therefore, because disorder acts locally, it does not remove the protected surface states. Secondly, several works used flux response and dislocation defects as a signature of higher order topology [11, 12, 13, 14, 15, 16, 17, 18], generalizing the Laughlin argument. This is reminiscent of the spectral flow in strong topological phases. We thus identify two open questions:
@@ 2 Why the extrinsic scattering invariant is not enough
The most direct approach to applying a scattering invariant to a HOTI is to use the appearance of the protected zero modes on the edge. This requires probing a reflection matrix of a finite symmetric sample with leads attached to its edges in a symmetric way. Symmetry class–dependent functions of the reflection matrix at each lead count the number of protected zero modes at the edge, and their parity was used as a scattering probe for HOTIs [3]. This extrinsic scattering invariant, and similar real space HOTI invariants [26, 27], rely on knowing the position of protected edge states, and therefore they are limited in usefulness for characterizing an unknown system.
@@ 3 Intrinsic scattering invariant construction
where $H$ is the Hamiltonian of the scattering region and the leads, and $\Psi^{\textrm{in}}$, $\Psi^{\textrm{out}}$, and $\Psi^{\textrm{localized}}$ are the wavefunctions of the incoming, outgoing, and localized states, respectively. These wavefunctions are matrices with columns corresponding to different modes, see App. A for a detailed description. To treat topology due to antisymmetries that map $E$ to $-E$ on the same footing, we set the energy $E$ to zero. The incoming and outgoing wavefunctions occupy the same lead sites and orbitals, while the localized wavefunctions are only defined in the scattering region. The total number of modes is the number of columns of the wavefunction matrices $N_{\textrm{modes}}=N_{\textrm{sites}}\times N_{\textrm{orbitals}}$, where $N_{\textrm{sites}}$ is the number of sites to which the leads are attached, and $N_{\textrm{orbitals}}$ is the number of orbitals per site in the lead. The scattering equations define a valid solution for any combination of amplitudes $\alpha^{\textrm{in}}$ of the incoming modes. Because the scattering matrix in Eq. (1) specifies a relation between the amplitudes of the incoming and outgoing wavefunctions, the scattering matrix is a linear map rather than an operator. Separating the leads into two groups imposes a block structure on the scattering matrix:
@@
Figure 3: Scattering geometries that probe the existence of zero-energy surface states using the bulk symmetries. The reflection matrix $r$ encodes the amplitudes of the wavefunctions that are reflected back into the lead (purple) from the scattering region (turquoise). (a) Scattering setup used to probe topological phases protected by local symmetries. A two-dimensional bulk Hamiltonian defines a one-dimensional reflection matrix by attaching leads along one direction and applying twisted boundary conditions in the other. The reflection matrices $r$ and $r^{\prime}$ capture the appearance of zero-energy bound states (b). (c) Scattering setup used to probe topological phases protected by spatial symmetries. A two-dimensional six-fold symmetric bulk Hamiltonian defines a one-dimensional reflection matrix by attaching symmetric leads and introducing a flux through the center of the scattering region. The blue area shows the matrix elements of the symmetry operator that gain a phase factor due to the flux, defining $C_{6}(\phi)$. The scattering setup probes the reflection from the outer lead, which is composed of six disjoint parts related to each other by the rotation symmetry.
The scattering theory of topological invariants [20] relates the topology of a $d$-dimensional bulk Hamiltonian $H_{\textrm{bulk}}(\bm{k}_{d})$, with $\bm{k}_{d}$ the wave vector, to the topology of a $d-1$-dimensional effective Hamiltonian $H_{\textrm{eff}}(\bm{k}_{d-1})$. The dimensional reduction procedure to obtain this effective Hamiltonian is as follows:
@@
    $H_{\textrm{eff}}=\begin{cases}r=r^{\dagger}&\text{if }H_{\textrm{bulk}}\text{ has chiral symmetry,}\\ \begin{pmatrix}0&r\\ r^{\dagger}&0\end{pmatrix}&\text{otherwise}.\end{cases}$ (3)
This effective Hamiltonian has zero eigenvalues simultaneously with $r(\bm{k}_{d-1})$ and a symmetry class different from the symmetry of $H_{\textrm{bulk}}$. The zero eigenvalues of $r(\bm{k}_{d-1})$ correspond to quantized transmission eigenvalues from one lead to the other, which only happens if the gap of $H_{\textrm{bulk}}$ closes. Because of this relation, the topological invariants of $H_{\textrm{eff}}$ and $H_{\textrm{bulk}}$ are equal. An alternative interpretation of the scattering invariant is to consider a boundary between the topological bulk and a trivial region. In this case, the boundary Green’s function defines the topology of the bulk [31]. In the scattering description this boundary corresponds to interrupting the lead with a reflection matrix $r^{\prime}$ of the trivial region, as shown in Fig. 3(b). The zero energy modes at an interface of an infinite system appear whenever $rr^{\prime}$ has an eigenvalue equal to $1$, and therefore the protected zero energy solutions at the interface are encoded in $r$.
While the procedure above may be applied to intrinsic HOTIs, the resulting scattering geometry either has leads swapped by the symmetry of the HOTI, or is incompatible with the global spatial symmetry protecting the HOTI phase. This incompatibility makes the topology of $H_{\textrm{eff}}$ different from the higher order topology of $H_{\textrm{bulk}}$. Instead, in order to apply the scattering approach to an intrinsic HOTI, the scattering geometry must be compatible with the global spatial symmetry protecting the HOTI phase, and this symmetry must map the leads onto themselves. At the same time, the two leads must be separated by the bulk rather than belonging to the same surface, such that no topological boundary modes propagate between them. This is required in order for zeros of the reflection matrix to correspond to the bulk gap closing and therefore to probe only bulk topological transitions. Otherwise, zeros of the reflection matrix could appear due to the closing of the surface gap and therefore not probe the bulk topology, as illustrated in Sec. 2. To satisfy the requirements of the scattering theory we propose to use a finite geometry with a hole: an annulus in two dimensions or a cylinder in three dimensions, and attach one lead to the inside and one lead to the outside of the sample. Figure 3(c) shows such a scattering geometry with both the inner and outer leads consisting of several disjoint parts that are related among themselves by the rotating symmetry. Whether the leads are attached to the entire circumference or only a part of it changes the magnitude of the finite size effects, but keeps the topological properties of the scattering geometry the same.
@@
Similar to the dimensional reduction procedure, we apply twisted boundary conditions in the direction parallel to the cylinder axis, which introduces a wave vector $k_{d-2}$. Without introducing additional parameters, this would result in a $d-2$-dimensional reflection matrix, which loses information about the angular direction along the cylinder and is unlikely to be sufficient to determine the HOTI topology. To preserve the information about the angular direction, we therefore introduce a flux $\phi$ through the center of the cylinder, which acts as an additional momentum parameter. The resulting procedure maps a $d$-dimensional bulk onto a $d-1$-dimensional $H_{\textrm{eff}}(\phi,\bm{k}_{d-2})$, similar to the dimensional reduction of the strong topological insulators. Contracting the inner radius of the cylinder to a point generally gaps out the surface modes due to finite size effects. Therefore, the flux is also necessary to cancel this splitting, as shown in Fig. 4. This cancellation of the finite size splitting was reported in Ref. [32], in a specific symmetry class. By confirming that the scattering invariants work in the annulus geometry, we show that the flux response is a general feature of the higher order topology.
@@
where $\tilde{H}(R_{\mathcal{O}}(\phi,\bm{k}_{d-2}))=\mathcal{O}H(\phi,\bm{k}_{d-2})\mathcal{O}^{-1}$ is the transformed Hamiltonian, and $R_{\mathcal{O}}$ is the action of the operator in parameter space, which in general depends on flux and momenta. To obtain the transformed scattering matrix $\tilde{S}(R_{\mathcal{O}}(\phi,\bm{k}_{d-2}))$, we act with the operator on $\mathcal{O}$ on the wavefunctions and coefficients of the wavefunctions, and equate the coefficients of the wavefunctions in Eq. (1) and Eq. (4) (see App. B for a detailed derivation). The action of an operator on the lead wavefunctions depends on whether this operator is unitary or antiunitary, and whether it is a symmetry or an antisymmetry of the Hamiltonian. We denote the unitary symmetries $\mathcal{U}$, unitary antisymmetries $\mathcal{C}$ (chiral symmetry), antiunitary symmetries $\mathcal{T}$ (time-reversal symmetry), and antiunitary antisymmetries $\mathcal{P}$ (particle-hole symmetry). The action of these 4 types of operators on the lead modes is [20] [^2] Note that we choose a different convention for $V$ and $Q$ than in Ref. [20]. Swapping $V\leftrightarrow V^{T}$ and $Q\leftrightarrow Q^{T}$ would make our convention equivalent to Ref. [20].:
@@
where $V_{\mathcal{O}}(\phi,\bm{k}_{d-2})$ and $Q_{\mathcal{O}}(\phi,\bm{k}_{d-2})$ are unitary matrices, which we compute by projecting the operator $\mathcal{O}$ onto the lead wavefunctions. By combining the transformed scattering equation Eq. (4) with the transformed wavefunctions in Eqs. (5), we obtain the transformed reflection matrix $\tilde{r}$:
@@
Due to $r(\phi,\bm{k}_{d-2})$ being a linear map (as a submatrix of $S$), the symmetry operations (6) apply to $r(\phi,\bm{k}_{d-2})$ differently than to $H_{\textrm{bulk}}$. Additionally, due to Eq. (3), whenever $H_{\textrm{bulk}}$ has chiral symmetry, $H_{\textrm{eff}}$ does not, and vice versa. This results in the Altland–Zirnbauer symmetry class of $H_{\textrm{bulk}}$ being related to that of $H_{\textrm{eff}}$ according to the Bott periodicity [20]. For example, in the symmetry class D ($\mathcal{P}^{2}=1$), particle-hole symmetry constraint is equivalent to $H_{\textrm{bulk}}(\mathbf{k})=-H_{\textrm{bulk}}^{*}(-\mathbf{k})$ up to a basis choice. The reflection matrix symmetry constraint (6b) is then equivalent to $r(-\phi,\bm{k}_{d-2})=r^{*}(-\phi,-\mathbf{k}_{d-2})$. After applying (3), the effective Hamiltonian belongs to the symmetry class BDI ($\mathcal{T}^{2}=\mathcal{P}^{2}=1$).
@@
In practice, the first step in determining the topological invariant requires finding a basis for the incoming and outgoing wavefunctions where the symmetry constraints on the reflection matrix are minimal. Choosing an appropriate basis simplifies the time-reversal symmetry constraint to either symmetry or antisymmetry of $r(\phi,\bm{k}_{d-2})$, particle-hole symmetry becomes a reality constraint, and chiral symmetry becomes a Hermiticity constraint, as shown in App. B. In the presence of multiple symmetries, we choose the basis where the chiral symmetry gives a Hermiticity constraint if it is present, according to Eq. (3), and the other symmetries have a minimal compatible form, following Ref. [20]. After performing this transformation, we rely on the established theory of Hamiltonian topological invariants to determine the topological invariant of $H_{\textrm{eff}}$.
@@ 4.1 Intrinsic scattering invariant of BBH model
To construct the intrinsic scattering invariant of the BBH model, we first consider the reflection matrix of an annulus-shaped lattice with a flux $\phi$ and inner and outer leads, as shown in Fig. 5(a). In the topological phase, both surfaces of the annulus host four zero-dimensional protected zero modes. Four-fold rotation and sublattice symmetries given by Eqs. (6a) and (6d) constrain the reflection matrix $r(\phi)$ as
$r(\phi)=Q_{C_{4}}(\phi)\>r(\phi)V_{C_{4}}^{\dagger}(\phi)=V_{\mathcal{C}}\>r^{\dagger}(\phi)Q_{\mathcal{C}}^{\dagger},$ (10)
@@
As the first step, we simplify the problem by observing that $C_{2}\equiv C_{4}^{2}$ commutes both with $\mathcal{C}$ and $\mathcal{C}_{4}$. Because $C_{2}^{2}(\phi)=-\exp(i\phi)$, interpreting $\phi$ as a momentum parameter of the effective Hamiltonian means that the problem is invariant under translation by half a unit cell in the corresponding direction. Therefore, we unfold the Brillouin zone and reduce the unit cell to a primitive one to eliminate redundant degrees of freedom. Because the eigenvalues of $C_{2}$ are ±$\pm i\exp(i\phi/2)$, they are periodic over $4\pi$ and the two subspaces of $C_{2}$—and of any operator that commutes with it—swap over a $2\pi$ interval. Therefore, we project $r$ and the remaining symmetries onto the subspace of $C_{2}$ with the eigenvalue $i\exp i\phi/2$, and consider their dependence on $\phi\in[0,4\pi)$. The range of $\phi\in[2\pi,4\pi)$ of the $i\exp i\phi/2$ eigenspace is equivalent to the range $\phi\in[0,2\pi)$ of the $-i\exp i\phi/2$ eigenspace, so choosing one subspace over twice the interval provides an equivalent representation. To avoid dealing with $4\pi$ periodicity, we redefine $\phi:=\phi/2$, which yields $r(\phi)$ that satisfies symmetry constraints equivalent to those in Eq. (10), but with $V_{C_{4}}^{2}=Q_{C_{4}}^{2}=i\exp i\phi$ instead. We call this procedure factoring out a symmetry, because $r(\phi)$ is no longer constrained by $C_{2}$. We describe how to factor out a symmetry and transform the remaining symmetries to the new basis in App. C.
@@
where the blocks of $H_{\textrm{eff}}(\phi)$ are in the basis of the different eigensubspaces of $V_{C_{4}}$. The eigensubspaces of $V_{C_{4}}$ swap under $\phi\to\phi+2\pi$, and therefore the effective Hamiltonian in Equation (13) is $4\pi$-periodic in $\phi$. Contrary to the dimensional reduction procedure used for strong topological insulators, this Hamiltonian has a sublattice-like symmetry—more specifically, a glide sublattice symmetry—just like its parent BBH model. The final step is to apply the topological invariant of $H_{\textrm{eff}}(\phi)$, however, to the best of our knowledge, the topology of one-dimensional systems combining fractional translations and sublattice symmetries has not been studied before. Therefore, we proceed to construct the invariant, although in most cases one may apply the already established invariants to the effective Hamiltonian.
Because $H_{\textrm{eff}}(\phi)$ is gapped as long as $h(\phi)$ has no zero eigenvalues, its topological invariant must be a function of $\det h(\phi)$. Because $\det h(\phi+2\pi)=\det h^{\ast}(\phi)$, $\det h(\phi)$ crosses the real axis an odd number of times in the interval $\phi\in[0,2\pi)$. The $\mathbb{Z}_{2}$ invariant of this system is the parity of the number of the crossings by $\det h(\phi)$ of the positive real axis. The crossings of the real axis appear and disappear only in pairs, hence the parity of the number of crossings changes only if $\det h(\phi)=0$ for some $\phi$. Therefore we identify the scattering invariant of the BBH model to be:
@@
We construct the model and compute the scattering matrix using Kwant [33]. To demonstrate that the scattering invariant applies to disordered systems, we add a $C_{4}$-symmetric disorder to the BBH model. Specifically, we choose the intra-cell hopping equal to $\lambda[1+\delta(x,y)]$, where $\delta(x,y)=\delta(y,-x)$ is symmetric under $C_{4}$ rotations, and each $\delta(x,y)$ not related by symmetry is an independent normally distributed random variable with zero mean and standard deviation $1/2$. Enforcing the global symmetry of the disorder distribution is necessary to compute the invariant, because we use the subspaces of the symmetry operator to project the reflection matrix. We present the resulting topological invariant for a single disorder realization in Fig. 5. The invariant correctly changes at the phase transition of the BBH model, even when the lead is attached far from the corner charges. Because scattering invariants rely on the unitarity of the reflection matrix, the phase transition point is sensitive to finite-size effects that may arise from the overlap of the wavefunctions between the inner and outer radius of the annulus. Figures 5(b-f) show the gap of the reflection matrix $\Delta=\min|\det r(\phi)|$ and the invariant $\mathcal{Q}$ as a function of $\lambda/\gamma$ for different outer radii $r$ of the annulus. In the thermodynamic limit of $r\to\infty$ and complete absence of disorder, the invariant changes sign at $\lambda=\gamma$, as expected from the bulk topological invariant of the BBH model. Regardless of the presence of disorder or the size of the annulus, the invariant is always quantized and changes sign only if the reflection matrix has a zero eigenvalue.
Figure 5: Two scattering geometries of the BBH model and their intrinsic invariant in the presence of globally symmetric disorder. (a/d) Annulus geometry with a flux $\phi$ threaded through the center and leads (purple) attached to the inner and outer rings. In (a) the leads are attached where the zero modes are located and in (d) the leads are in between the zero modes. The gap of the reflection matrix $\Delta=\min|\det r(\phi)|$ in (b/e) closes simultaneously with the sign change of the invariant $\mathcal{Q}$ in (c/f). The results are shown for different outer ring sizes $r=10,20,40$, and in (c/f) the values of $\mathcal{Q}$ are minimally shifted for clarity.